How would you solve the equation
L=2.43*W^0.3326 if w=3

Answer by stanbon(75887)

How would you solve the equation
L=2.43*W^0.3326 if w=3
---
Use a calculator to get:
L = 2.43*3^0.3326 = 3.5018
=========================
Cheers,
Stan H.